# The Boundaries of Tractability in Hierarchical Task Network Planning

### v1

@@ Abstract
We study the complexity-theoretic boundaries of tractability for three classical problems in the context of Hierarchical Task Network Planning: the validation of a provided plan, whether an executable plan exists, and whether a given state can be reached by some plan. We show that all three problems can be solved in polynomial time on primitive task networks of constant partial order width (and a generalization thereof), whereas for the latter two problems this holds only under a provably necessary restriction to the state space. Next, we obtain an algorithmic meta-theorem along with corresponding lower bounds to identify tight conditions under which general polynomial-time solvability results can be lifted from primitive to general task networks. Finally, we enrich our investigation by analyzing the parameterized complexity of the three considered problems, and show that (1) fixed-parameter tractability for all three problems can be achieved by replacing the partial order width with the vertex cover number of the network as the parameter, and (2) other classical graph-theoretic parameters of the network (including treewidth, treedepth, and the aforementioned partial order width) do not yield fixed-parameter tractability for any of the three problems.
@@ 1 Introduction
Automated Planning is a core research topic whose goal is to devise computational methods capable of finding solutions to prominent planning tasks. In this work, we consider automated planning in the context of Hierarchical Task Networks (HTNs), which have received significant attention in the artificial intelligence research community. HTNs are capable of incorporating the compound, hierarchical structure of real-world tasks, as opposed to the fine-grained view of classical planning that operates directly on the state space of the system under consideration, and are one of the best-established planning formalisms in the literature Tsuneto et al. (1996); Kuter et al. (2005); Li et al. (2009); Hogg et al. (2009); Parunak et al. (2009); Sohrabi et al. (2009); Geier and Bercher (2011); Alford et al. (2015); Xiao et al. (2017); Höller et al. (2019); Behnke et al. (2019); Olz et al. (2021); Lin and Bercher (2023).
@@
On a high level, [^1] Formal definitions are provided in the Preliminaries. HTNs can be informally described and illustrated as follows: at the top level of the task hierarchy is a sequence of tasks that is to be executed in a (partially specified) order. For instance, the task of going to work may be decomposed into three subtasks: leaving the house, going the actual distance, and getting set up at the office. Now, each of these tasks can in turn be decomposed into a set of smaller tasks, like climbing the stairs or taking a bus, which again can be decomposed, and so forth. This continues until we reach an elementary level of tasks that are no longer decomposable, which correspond to single executable actions.
@@
The crucial feature of the model is that it separates the user-facing description of a task (what we want to do) from its internal implementation (how to do it). For instance, leaving the house always needs to be done before going to work, regardless of how it is done. However, in an HTN, we may have several options of how to decompose this task: the agent can, e.g., either climb the stairs or take the elevator. Similarly, the agent might opt for walking to work instead of going by bus.
In the literature, HTN planning is usually considered through the lens of heuristic approaches, which are evaluated against real-world problems. On the other hand, much less is known about the complexity-theoretic foundations of planning problems associated with HTNs. This lack of understanding contrasts with the major advances that have been made towards obtaining a thorough comprehension of the complexity of other fundamental problems arising in the context of artificial intelligence, such as Bayesian Network Learning Ordyniak and Szeider (2013); Grüttemeier et al. (2021a, b); Grüttemeier and Komusiewicz (2022), Clustering Ganian et al. (2020b, 2022); Kellerhals et al. (2023); Bandyapadhyay et al. (2023), Resource Allocation Bliem et al. (2016); Deligkas et al. (2022); Eiben et al. (2023), and Scheduling Ganian et al. (2020a); Bentert et al. (2023); Heeger et al. (2023). The aim of this article is to fill this gap by identifying the precise boundaries of tractability for the central computational tasks that arise in the context of HTN planning.
Considered Problems.
Perhaps the most natural algorithmic question associated with HTNs is the following: given an HTN, an initial state, and a plan, is the plan executable on the HTN? This question has been extensively investigated in the literature, where authors typically consider a plan to be formalized as either a sequence of tasks Behnke et al. (2015); Barták et al. (2021), or actions Höller et al. (2022); Lin and Bercher (2023). Here, we adopt the latter approach and note that the complexity of this problem—denoted Plan Validation—has been studied, e.g., in the context of post-optimization and repairing of solutions to HTN problems  Lin and Bercher (2023).
The second problem we consider is Plan Existence: given an HTN and some initial state, is there a way to execute all the tasks in the network? The complexity of this fundamental problem has been studied in several variants of HTN planning problems Erol et al. (1996); Geier and Bercher (2011); Alford et al. (2015); Chen and Bercher (2021).
Third, we consider the question of whether a specified target state can be reached from an initial state of the system, formalized as the State Reachability problem. This is the typical goal in classical planning Fikes and Nilsson (1971); Chapman (1987); Bäckström and Nebel (1995), so it should come as no surprise that HTN planning problems with specified target states have attracted interest as well Höller and Bercher (2021); Olz et al. (2021).
Motivated by the above considerations of states and tasks, we also define the corresponding problem of constructing a plan executing (at least) a specified set of actions. However, this Action Executability problem is in some sense a generalization of Plan Existence, and, in fact, all results that we prove for one of the problems also hold for the other. Thus, we omit Action Executability from the discussion of our results below.
Our Contribution.
The lack of results on the computational complexity of Plan Validation, Plan Existence, and State Reachability may at least partially stem from the fact that they remain computationally intractable (in particular, NP-hard) even in severely restricted settings. While the emerging intractability might seem difficult to navigate, we identify interesting and natural special cases where pockets of polynomial-time solvability show themselves. Moreover, we prove that these pockets give rise to a surprisingly rich complexity-theoretic landscape.
@@
To be more precise, we embark by considering primitive task networks, i.e., HTNs in which there are no compound tasks. While this may seem restrictive, it is a natural place to start from a complexity-theoretic perspective: all three of the considered problems remain NP-hard on primitive task networks. In fact, as our first results, we establish the NP-hardness of Plan Validation, Plan Existence, and State Reachability even on networks which consist of collections of stars (Theorems 5 and 2). Moreover, unlike Plan Validation, we show that the latter two problems remain NP-hard even on trivial (in particular, edgeless) networks when the state space is not of fixed size (Theorem 1); this means that in order to avoid immediate intractability, we henceforth consider Plan Existence and State Reachability exclusively in the setting of constant-sized state spaces. This requirement matches analogous restrictions that were used in other planning contexts Kronegger et al. (2013); Bäckström et al. (2015).
While these initial results may seem discouraging, we prove that all three problems are polynomial-time solvable on networks which are either chains or antichains (Observations 6 and 7, Theorems 10 and 8). These tractable cases hint at cracks in the wall of computational intractability. We pry these cracks open by showing that these two restrictions actually represent special cases in a spectrum of tractable special cases: Plan Validation, Plan Existence, and State Reachability are polynomial-time solvable on networks of bounded “generalized partial order width”, that is, the partial order width of the network when ignoring isolated elements (Theorems 11, 13 and 14). This non-trivial result, established via a combination of branching techniques and deep insights into the behavior of optimal plans in the state space, unifies and generalizes the aforementioned tractability results on chains and antichains.
Building on the above results for the primitive case, in the second part of the paper, we establish an algorithmic meta-theorem that allows us to precisely describe the circumstances under which general tractability results for primitive task networks can be lifted to the compound case. Essentially, if
  1. 1.
    the total number of compound tasks in the network,
  2. 2.
    the maximum size of a network that a task can be decomposed into, and
  3. 3.
    the maximum depth of a decomposition
are all bounded by some constant, then any polynomial-time solvability result for any “reasonable” HTN planning problem on primitive task networks can be lifted to compound networks (Theorem 16). Crucially, we prove that this result is tight in the sense that tractability cannot be guaranteed if any of the three restrictions are dropped (Theorems 17, 19, 20 and 18).
@@
In the third and final part of the paper, we push beyond classical complexity and analyze Plan Validation, Plan Existence, and State Reachability from the perspective of the more fine-grained parameterized complexity paradigm Downey and Fellows (2013); Cygan et al. (2015). In parameterized complexity, one considers the tractability of problems not only w.r.t. the input size $n$, but also w.r.t. a specified numerical parameter $k$, where the aim is to obtain algorithms which run in time at most $f(k)\cdot n^{\mathcal{O}(1)}$ (so-called fixed-parameter algorithms); problems admitting such algorithms are called fixed-parameter tractable.
@@
The natural question from the parameterized perspective is whether the aforementioned tractability results for HTNs with constant-sized structural measures (Theorems 11, 13 and 14) can be lifted to fixed-parameter tractability when parameterized by the same measures. While analogous questions have been thoroughly explored—and often answered in the affirmative—in many other areas of artificial intelligence research Ordyniak and Szeider (2013); Eiben et al. (2021); Ganian and Korchemna (2021); Froese et al. (2022); Heeger et al. (2023), here we prove that the generalized partial order width does not yield fixed-parameter tractability for any of our problems on primitive task networks (Corollaries 21 and 22), at least under well-established complexity assumptions. This means that different structural parameterizations are required to obtain fixed-parameter tractability.
Further, Theorems 5 and 2 can already be seen to rule out fixed-parameter tractability for the considered problems on primitive task networks w.r.t. classical graph-theoretic measures such as treewidth Robertson and Seymour (1986), treedepth Nesetril and de Mendez (2012), and even the feedback edge number Uhlmann and Weller (2013); Ganian and Korchemna (2021). However, fixed-parameter tractability can still be achieved under the right parameterization: in Theorems 23 and 24, we obtain highly non-trivial fixed-parameter algorithms for Plan Validation, Plan Existence, and State Reachability when parameterized by the vertex cover number of the input primitive task network. Finally, for compound HTNs, we lift our general algorithmic meta-theorem for polynomial-time solvability (Theorem 16) to the setting of fixed-parameter tractability when additionally parameterizing by the “breadth” of an HTN, i.e., the maximum number of pairwise non-isomorphic networks that a task can be decomposed into (Theorem 25).
@@
HTN planning has been extensively studied in the artificial intelligence research community and has found applications in areas such as healthcare García-Remesal (2008); González-Ferrer et al. (2011, 2013b); Bihlmaier et al. (2014), business processes González-Ferrer et al. (2013a); Ko et al. (2009, 2012), human-robot interaction Hayes and Scassellati (2016); Tewari and Persiani (2021), real-time games Ontañón and Buro (2015); Namiki et al. (2021); Soemers and Winands (2016), emergency response Liu et al. (2016); Zhao et al. (2017), and visualization Padia et al. (2019). For an overview of the current state of hierarchical planning at large, we refer the reader to the survey of Bercher, Alford, and Hoeller (2019) and the overview article of Georgievski and Aiello (2015).
The works most closely related to ours are those of Lin and Bercher (2023) and Olz, Biundo, and Bercher (2021), which also established some of the formalisms that we build on. In a similar vein, questions on the complexity and even computability of many tasks associated with HTNs have been treated in the literature Geier and Bercher (2011); Alford et al. (2015), where precursors of the abovementioned formal models were defined. Beyond these recent works, there are also classical investigations of computational questions related to HTN planning in slightly different settings than the ones we look at here, such as the early influential works by Erol, Hendler, and Nau (1994; 1996).
Albeit carried out in a different context, the work of Kronegger, Pfandler, and Pichler (2013) on propositional strips with negation (PSN) studies a problem ($k$-PSN) which is related to State Reachability. The main difference is that while PSN includes a broader definition of actions, it does not consider any primitive or hierarchical network structure whatsoever. Finally, we note that the computational complexity of planning has also been studied in a variety of other settings beyond HTNs Bylander (1994); Bäckström et al. (2012, 2013); Bäckström et al. (2015).
@@ 2 Preliminaries
For a positive integer $n$, let $[n]={\{1,\ldots,n\}}$. We employ standard graph-theoretic terminology Diestel (2012). The underlying undirected graph of a digraph $D$ is the graph obtained by replacing each arc in $D$ with an edge (discarding loops and multiple edges).
For some $V^{\prime}\subseteq V$, the subgraph induced by $V^{\prime}$ is $G[V^{\prime}]=(V^{\prime},E\cap(V^{\prime}\times V^{\prime}))$. A walk in $G$ is a sequence of vertices $v_{1},\ldots,v_{\ell}$ such that $(v_{i},v_{i+1})\in E$ for all $i\in[\ell-1]$. As the walk contains $\ell-1$ edges, we say that it has length $\ell-1$. A walk is a cycle if $v_{1}=v_{\ell}$. A graph is acyclic if it does not contain any cycles. A (simple) path is a walk such that $v_{i}\neq v_{j}$ for all $i\neq j,i,j\in[\ell]$. A simple cycle is a walk where this restrictions holds except for $v_{1}$ and $v_{\ell}$.
A multigraph is a graph that allows for multiple edges between a pair of vertices. Formally, we define a multigraph by $G=(V,E,\beta\colon E\rightarrow(V\times V))$. A vertex walk in $G$ is a sequence of vertices $v_{1},\ldots,v_{\ell}$ such that there exists $e_{i}\in E$ with $\beta(e_{i})=(v_{i},v_{i+1})$ for all $i\in[\ell-1]$. An edge walk in $G$ is a sequence of edges $e_{1},\ldots,e_{\ell}$ such that $e_{i}\in E$ for all $i\in[\ell]$ and if $\beta(e_{i})=(v,v^{\prime})$, then $\beta(e_{i+1})=(v^{\prime},v^{\prime\prime})$ for some $v^{\prime\prime}\in V$. Note that an edge walk corresponds to a vertex walk by considering the vertices it passes through. We transfer the notion of (simple) vertex paths and cycles and (simple) edge paths and cycles from non-multigraphs, where an edge walk is simple if and only if its corresponding vertex walk is simple.
The partial order width $w$ of a partially ordered set $\lesssim$ is the maximum number of pairwise-disjoint maximal chains contained in $\lesssim$ or, equivalently, the size of its largest antichain Dilworth (1950). An element in $\lesssim$ is isolated if it is incomparable to every other element in the set. We define the generalized partial order width of a partially ordered set $\lesssim$ as the maximum number of pairwise-disjoint maximal chains contained in $\lesssim$ which are not isolated elements (or, equivalently, the size of the largest antichain not containing any isolated element).
@@
Our formalization of HTNs follows the terminology employed in recent works Olz et al. (2021); Lin and Bercher (2023). The domain of a planning problem is a tuple $(F,A,{\cal C},\delta,M)$, where $F$ is a finite set of propositions, $A$ is a set of actions (or primitive task names), and ${\cal C}$ is a set of compound task names such that $A\cap{\cal C}=\emptyset$. The function $\delta\colon A\rightarrow 2^{F}\times 2^{F}\times 2^{F}$ maps each action to a set $\operatorname{prec}(a)$ of preconditions, a set $\operatorname{del}(a)$ of propositions to be removed, and a set $\operatorname{add}(a)$ of propositions to be added to the current state. Lastly, $M$ represents the decomposition methods and is a mapping from each $c\in{\cal C}$ to a set $Z$ of task networks (defined below) which $c$ can be decomposed into.
A task network in the domain $(F,A,{\cal C},\delta,M)$ is defined as a tuple $(T,\prec^{+},\alpha)$ that consists of a set $T$ of task identifiers with a partial order $\prec^{+}\ \subseteq\colon T\times T$ and a function $\alpha\colon T\rightarrow A\cup{\cal C}$. The usual representation considered for a task network is its cover graph $D_{\prec}=(T,\prec)$, where $\prec$ is the cover of $\prec^{+}$ (in particular, $(a,b)\in\prec$ if and only if $a\prec^{+}b$ and there is no element $c$ such that $a\prec^{+}c\prec^{+}b$). When analyzing the structure of directed graphs, it is often useful to also consider their underlying undirected graph Ordyniak and Szeider (2013); Ganian et al. (2021); we hence use $G_{\prec}$ to denote the underlying undirected graph of $D_{\prec}$. An example of a simple task network is provided in Fig. 1.
@@
Figure 1: The digraph $D_{\prec}$ for a primitive HTN with tasks $T={\{t_{1},t_{2},t_{3},t_{4}\}}$, actions $A={\{a_{1},a_{2},a_{3}\}}$, propositions $F=\{1,2\}$, and initial state $s_{0}=\emptyset$. Each task’s action is given in its node. A directed edge from a task $t$ to a task $t^{\prime}$ corresponds to $(t,t^{\prime})\in\prec^{+}$.
A set of tasks $T^{\prime}\subseteq T$ is a chain if $\prec^{+}$ defines a total order on $T^{\prime}$, and an antichain if all tasks in $T^{\prime}$ are pairwise incomparable w.r.t. $\prec^{+}$. Two task networks $tn=(T,\prec^{+},\alpha)$ and $tn^{\prime}=(T^{\prime},{\prec^{+}}^{\prime},\alpha^{\prime})$ are isomorphic if there is a bijection $\phi\colon T\rightarrow T^{\prime}$ such that $(t,t^{\prime})\in\prec^{+}$ if and only if $(\phi(t),\phi(t^{\prime}))\in{\prec^{+}}^{\prime}$ and $\alpha(t)=\alpha^{\prime}(\phi(t))$ for all $t,t^{\prime}\in T$. Intuitively, two task networks are isomorphic if they describe the same structure, but using different task names.
A task $t$ is primitive if $\alpha(t)\in A$ and compound otherwise. A task network is primitive if it contains only primitive tasks. Non-primitive task networks can be decomposed into primitive ones by applying the decomposition methods in $M$. More precisely, let there be a task network $tn=(T,\prec^{+},\alpha)$, $t\in T$, and $m=(\alpha(t),tn_{m})\in M$. Let $tn_{m}^{\prime}=(T_{m}^{\prime},{\prec_{m}^{+}}^{\prime},\alpha_{m}^{\prime})$ with $T_{m}^{\prime}\cap T=\emptyset$ being isomorphic to $tn_{m}$. Then, $t$ in $tn$ can be decomposed by $m$ to obtain the new task network $tn^{\prime}=(T^{\prime},{\prec^{+}}^{\prime},\alpha^{\prime})$ with
$T^{\prime}$ $=T\setminus{\{t\}}\cup T_{m}^{\prime},$
${\prec^{+}}^{\prime}$ $=\prec^{+}\cap(T^{\prime}\times T^{\prime})\cup{\prec^{+}_{m}}^{\prime}\cup\{(t_{1},t_{2})\mid(t_{1},t)\in\prec^{+},$
$t_{2}\in T_{m}^{\prime}\}\cup{\{(t_{2},t_{1})\mid(t,t_{1})\in\prec^{+},t_{2}\in T_{m}^{\prime}\}},$
$\alpha^{\prime}$ $=\alpha\setminus{\{(t,\alpha(t))\}}\cup\alpha_{m}^{\prime}.$
Intuitively, decomposing a task $t$ to a subnetwork $tn_{m}$ means replacing $t$ by $tn_{m}$ such that each task in $tn_{m}$ is connected to tasks outside $tn_{m}$ in the same way $t$ was before. The extra step using the isomorphic $tn_{m}^{\prime}$ is required to ensure unique task identifiers after the decomposition. An example is provided in Fig. 2.
Figure 2: Example for the decomposition of a compound task network. Let $tn$ be the compound task network that is depicted in Fig. 1, where $t_{3}$ is a compound task instead of a primitive task with action $a_{1}$. If $(\alpha(t_{3}),tn_{m})\in M$, then $tn$ decomposes to $tn^{\prime}$.
An HTN planning problem is a problem whose input includes a domain $(F,A,{\cal C},\delta,M)$, an initial task network $tn$ from that domain, and an initial state $s_{0}\subseteq F$. An action $a$ is executable in a state $s\subseteq F$ if $\operatorname{prec}(a)\subseteq s$. If executed, it changes the state $s$ to $s^{\prime}=(s\setminus\operatorname{del}(a))\cup\operatorname{add}(a)$. A plan is a sequence of actions. A plan is executable from an initial state $s_{0}$ if all actions in the plan can be executed in sequence (one after the other) starting from $s_{0}$. A sequence $t_{1},t_{2},\ldots,t_{\ell}$ of tasks is executable in a task network $tn$ if $tn$ can be decomposed into a primitive task network $tn^{\prime}$ such that the plan $\alpha(t_{1}),\ldots,\alpha(t_{\ell})$ is executable, $t_{i}\neq t_{j}$, and $t_{i}\not\prec^{+}t_{j}$ if $i>j$. For a sequence $p$ of tasks or actions, we refer to its $i^{\text{th}}$ element by $p[i]$ and to its length by $|p|$.
Example 1.
Consider the primitive HTN depicted in Fig. 1. The task sequence $t_{1},t_{2},t_{4}$ is not executable since $(t_{3},t_{4})\in\prec^{+}$ means that $t_{3}$ must be executed before $t_{4}$. The task sequence $t_{1},t_{3}$, $t_{4}$ is not executable either, as after executing $t_{1}$ and $t_{3}$, the state is ${\{1\}}$, so the precondition of $a_{3}$ is not fulfilled when executing $t_{4}$. On the other hand, the task sequence $t_{1},t_{2},t_{3},t_{4}$ is executable. It conforms with the partial order, and for the corresponding action sequence $a_{1},a_{2},a_{1},a_{3}$, after executing any prefix of the sequence, the state is a superset of the precondition of the next action.
We are now ready to define our problems of interest.
Plan Validation
@@
Input: A task network $tn=(T,\prec^{+},\alpha)$ over some domain $(F,A,{\cal C},\delta,M)$, an initial state $s_{0}\subseteq F$, and a plan $p$ over $A$.
Question: Is there a task sequence that corresponds to plan $p$, is executable in $tn$ from $s_{0}$, and executes all the tasks in $tn$?
As mentioned in the introduction, authors have also sometimes considered a version of Plan Validation where we are provided a sequence of tasks on the input as opposed to a plan. We remark that this would give rise to a problem which is trivially linear-time solvable on all primitive task networks, and is also fully captured by our algorithmic meta-theorems (Theorems 16 and 25) on compound networks. In view of this, we exclude this problem from our further considerations.

### v2

# A Structural Complexity Analysis of Hierarchical Task Network Planning

@@ Abstract
We perform a refined complexity-theoretic analysis of three classical problems in the context of Hierarchical Task Network Planning: the verification of a provided plan, whether an executable plan exists, and whether a given state can be reached. Our focus lies on identifying structural properties which yield tractability. We obtain new polynomial algorithms for all three problems on a natural class of primitive networks, along with corresponding lower bounds. We also obtain an algorithmic meta-theorem for lifting polynomial-time solvability from primitive to general task networks, and prove that its preconditions are tight. Finally, we analyze the parameterized complexity of the three problems. In line with IJCAI standards and guidelines, full details for some proofs are deferred to the technical appendix.
@@ 1 Introduction
Automated Planning is a core research topic whose goal is to devise computational methods capable of finding solutions to prominent planning tasks. In this work, we consider automated planning in the context of Hierarchical Task Networks (HTNs), which have received significant attention in the artificial intelligence research community. HTNs are capable of incorporating the compound, hierarchical structure of real-world tasks, as opposed to the fine-grained view of classical planning that operates directly on the state space of the system under consideration, and are one of the best-established planning formalisms in the literature Tsuneto et al. (1996); Li et al. (2009); Sohrabi et al. (2009); Geier and Bercher (2011); Alford et al. (2015b); Xiao et al. (2017); Behnke et al. (2019); Höller et al. (2019); Lin and Bercher (2023).
@@
On a high level, HTNs can be informally described and illustrated as follows: at the top level of the task hierarchy is a set of tasks that is to be executed in a (partially specified) order. For instance, the task of going to work may be decomposed into three subtasks: leaving the house, going the actual distance, and getting set up at the office. Now, each of these tasks can in turn be decomposed into a set of smaller tasks, like climbing the stairs or taking a bus, which again can be decomposed, and so forth. This continues until we reach an elementary level of tasks that are no longer decomposable, which correspond to single executable actions.
@@
The crucial feature of the model is that it separates the user-facing description of a task (what we want to do) from its internal implementation (how to do it). For instance, leaving the house needs to be done before going the actual distance, regardless of how it is done. However, in an HTN, we may have several options of how to decompose this task: the agent can, e.g., either climb the stairs or take the elevator.
Perhaps the most natural problem associated with HTNs is Plan Verification: can the tasks in a given network be decomposed and then ordered to match a given, executable sequence of actions? A second fundamental problem is Plan Existence: can the tasks in a given network be decomposed and then ordered in any way that yields an executable sequence of actions? The literature has investigated the complexity of several variants of the aforementioned problems Erol et al. (1996); Geier and Bercher (2011); Alford et al. (2015b); Behnke et al. (2015); Lin and Bercher (2023). Moreover, we also consider the question of whether a specified target state can be reached, formalized as the State Reachability problem. This is the typical goal in classical planning Fikes and Nilsson (1971); Chapman (1987); Bäckström and Nebel (1995), and has also been considered in the HTN setting Höller and Bercher (2021); Olz et al. (2021).
It is well-established that problems on HTNs such as those listed above remain computationally intractable even in severely restricted settings. These results usually depend on restricting the decompositions, e.g., requiring acyclicity Alford et al. (2015a); Chen and Bercher (2021), restricting the impact of actions on the state space (e.g., delete-relaxed HTN planning) Alford et al. (2014); Höller et al. (2020), or by allowing the insertion of tasks at arbitrary positions Geier and Bercher (2011); Alford et al. (2015b).
In this article, we take a different approach and instead focus on structural restrictions of the task network. While there are some complexity results known for the extreme cases of a total ordering or no ordering of the tasks Geier and Bercher (2011); Alford et al. (2014); Behnke and Speck (2021), other restrictions received little to no attention so far. Building on these first results, we identify interesting and natural special cases where pockets of polynomial-time solvability show themselves. Moreover, we prove that these pockets give rise to a surprisingly rich complexity-theoretic landscape.
Our Contributions.
@@
We first consider primitive task networks, i.e., HTNs in which there are no compound tasks. While this may seem restrictive, it is a natural place to start from a complexity-theoretic perspective: all three of the considered problems remain NP-hard on primitive task networks. In fact, Plan Verification is known to be NP-hard even if the network forms a collection of disjoint stars Lin and Bercher (2023). Moreover, unlike Plan Verification, the latter two problems remain NP-hard even on trivial (in particular, edgeless) networks when the state space is not of fixed size Bylander (1994); Erol et al. (1996). Thus, in line with previous related works, we henceforth consider Plan Existence and State Reachability exclusively in the setting of constant-sized state spaces Kronegger et al. (2013); Bäckström et al. (2015). Even in this setting, we establish the NP-hardness of the three considered problems on networks consisting of a collection of independent chains.
While these initial results may seem discouraging, it turns out that all three problems are polynomial-time solvable on networks which are chains (i.e., totally-ordered networks) Erol et al. (1996); Behnke et al. (2015) or antichains (Theorems 3 and 4). We generalize and unify these tractability results by showing that all three considered problems are polynomial-time solvable on networks of bounded “generalized partial order width”, that is, the partial order width of the network when ignoring isolated elements (Theorems 5 and 6).
@@
Building on the above results for the primitive case, in the second part of the paper, we establish an algorithmic meta-theorem that allows us to precisely describe the conditions under which polynomial-time tractability for primitive task networks can be lifted to the compound case (Theorem 8). Crucially, we prove that this result is tight in the sense that all of the required conditions are necessary (Theorem 9). In the final part of the paper, we push beyond classical complexity and analyze the considered problems from the perspective of the more fine-grained parameterized complexity paradigm Cygan et al. (2015). There, one analyzes problems not only w.r.t. the input size $n$, but also w.r.t. a specified parameter $k\in\mathbb{N}$, where the aim is to obtain algorithms which run in time at most $f(k)\cdot n^{\mathcal{O}(1)}$ for some computable function $f$ (so-called fixed-parameter algorithms); problems admitting such algorithms are called fixed-parameter tractable (FPT).
@@
The natural question from the parameterized perspective is whether the aforementioned tractability results for HTNs with constant-sized structural measures (Theorems 5 and 6) can be lifted to fixed-parameter tractability when parameterized by the same measures. While analogous questions have been thoroughly explored—and often answered in the affirmative—in many other areas of artificial intelligence Eiben et al. (2021); Ganian and Korchemna (2021); Froese et al. (2022); Heeger et al. (2023), here we prove that the generalized partial order width does not yield fixed-parameter tractability for any of our problems on primitive task networks, under well-established complexity assumptions. Moreover, the aforementioned hardness of the problems on forests of stars already rules out fixed-parameter tractability for essentially all standard structural graph parameters, except for the vertex cover number (vcn) Blazej et al. (2023); Bodlaender et al. (2023); Chalopin et al. (2024). We complement these lower bounds by designing highly non-trivial fixed-parameter algorithms for all three considered problems on primitive task networks w.r.t. the vertex cover number of the network. For compound HTNs, we then adapt our general algorithmic meta-theorem for polynomial-time solvability to the fixed-parameter paradigm (Theorem 12), allowing us to also extend these results to the non-primitive setting. See Fig. 1 for a schematic overview of our results.
Hardness on Tree-like Graphs (1, 2) Tractability on (Anti)-Chains (3, 4) P for fixed gpow (5,6) Meta-Theorem (8) Lower Bound for Meta-Theorem (9) FPT w.r.t. vcn (10, 11) Algorithms for Compound Networks FPT Meta-Theorem (12) motivatemotivates
Figure 1: Overview of our results and respective theorems. Generalized Partial Order Width is abbreviated to gpow, and Vertex Cover Number to vcn. Results in gray boxes refer to primitive instances.
@@
HTN planning has been extensively studied in artificial intelligence research and has applications in areas such as healthcare González-Ferrer et al. (2013b), business processes González-Ferrer et al. (2013a), human-robot interaction Hayes and Scassellati (2016), emergency response Liu et al. (2016); Zhao et al. (2017), and visualization Padia et al. (2019). For an overview of the current state of HTN planning, see the works by Bercher, Alford, and Hoeller Bercher et al. (2019) and Georgievski and Aiello Georgievski and Aiello (2015).
Questions on the complexity and even computability of many tasks associated with HTNs have been treated in the literature Erol et al. (1994, 1996); Geier and Bercher (2011); Alford et al. (2015a); Olz et al. (2021); Lin and Bercher (2023); Lin et al. (2024a, b). Albeit carried out in a different context, the work of Kronegger, Pfandler, and Pichler Kronegger et al. (2013) on propositional STRIPS with negation (PSN) studies a problem ($k$-PSN) which is related to State Reachability. The main difference is that while PSN includes a broader definition of actions, it does not consider any primitive or hierarchical network structure. Finally, the computational complexity of planning has also been studied in settings beyond HTNs Bylander (1994); Bäckström et al. (2012, 2013); Bäckström et al. (2015).
@@ 2 Preliminaries
For $n\in\mathbb{N}$, let $[n]={\{1,\ldots,n\}}$. We use standard graph terminology Diestel (2012). The partial order width $w$ of a partially ordered set $\lesssim$ is the maximum number of pairwise-disjoint maximal chains contained in $\lesssim$ or, equivalently, the size of its largest antichain. An element in $\lesssim$ is isolated if it is incomparable to each other element in the set. We define the generalized partial order width of a partially ordered set $\lesssim$ as the partial order width of $\lesssim$ after removing isolated elements.
@@
Our formalization of HTNs follows the terminology employed in recent works Olz et al. (2021); Lin and Bercher (2023). The domain of a planning problem is a tuple $(F,A,{\cal C},\delta,M)$, where $F$ is a finite set of propositions, $A$ is a set of actions (or primitive task names), and ${\cal C}$ is a set of compound task names such that $A\cap{\cal C}=\emptyset$. The function $\delta\colon A\rightarrow 2^{F}\times 2^{F}\times 2^{F}$ maps each action to a set $\operatorname{prec}(a)$ of preconditions, a set $\operatorname{del}(a)$ of propositions to be removed, and a set $\operatorname{add}(a)$ of propositions to be added to the current state. An action $a$ is executable in a state $s\subseteq F$ if $\operatorname{prec}(a)\subseteq s$. If executed, it changes the state $s$ to $s^{\prime}=(s\setminus\operatorname{del}(a))\cup\operatorname{add}(a)$. A plan is a sequence of actions and is executable in a state $s$ if its actions can be executed consecutively starting from $s$. Lastly, $M$ represents the decomposition methods and maps each $c\in{\cal C}$ to a set of task networks (defined below) which $c$ can be decomposed into.
A task network in such a domain is a tuple $(T,\prec^{+},\alpha)$ consisting of a set $T$ of task identifiers (or simply tasks) with a partial order $\prec^{+}\ \subseteq T\times T$ and a function $\alpha\colon T\rightarrow A\cup{\cal C}$.
The usual representation of a task network is its cover graph $D_{\prec}=(T,\prec)$, where $\prec$ is the cover of $\prec^{+}$ (in particular, $(a,b)\in\prec$ if and only if $a\prec^{+}b$ and there is no $c$ such that $a\prec^{+}c\prec^{+}b$). We denote the underlying undirected graph by $G_{\prec}$. Fig. 2 provides an example of a task network.
@@
Figure 2: The digraph $D_{\prec}$ for a primitive HTN with tasks $T={\{t_{1},t_{2},t_{3},t_{4}\}}$, actions $A={\{a_{1},a_{2},a_{3}\}}$, propositions $F=\{1,2\}$, and initial state $s_{0}=\emptyset$. A task’s action is in its node. There is an arc from a task $t$ to a task $t^{\prime}$ if $(t,t^{\prime})\in\prec$. Here, the only sequences that execute all tasks are $t_{1},t_{2},t_{3},t_{4}$ and $t_{1},t_{3},t_{2},t_{4}$.
A subset $T^{\prime}$ of tasks is a chain if $\prec^{+}$ defines a total order on $T^{\prime}$, and an antichain if all of its tasks are pairwise incomparable. A task $t$ is primitive if $\alpha(t)\in A$, and compound otherwise. A task network is primitive if it contains no compound tasks. For a primitive task network, a total order $\bar{t}$ on $T$ respecting $\prec$ is called a linearization and we denote the corresponding action sequence by $\alpha(\bar{t})$. Non-primitive task networks can be decomposed into primitive ones via the methods in $M$; decomposing a task $t$ into a subnetwork $tn_{m}\in M(t)$ means replacing $t$ by $tn_{m}$ such that each task in $tn_{m}$ is connected to tasks outside $tn_{m}$ in the same way that $t$ was.
The input of an HTN problem contains a domain $D=(F,A,{\cal C},\delta,M)$, an initial task network $tn$ in that domain, and an initial state $s_{0}\subseteq F$111 Some works refer to the triple $(D,tn,s_{0})$ as an HTN problem, while we use that term for the considered algorithmic problems. In a generalization of the literature, we consider a primitive task network $tn^{\prime}=(T^{\prime},{\prec^{+}}^{\prime},\alpha^{\prime})$ to be a solution to a triple $(D,tn,s_{0})$ if $tn^{\prime}$ has a linearization $\bar{t}$ such that $\alpha^{\prime}(\bar{t})$ is executable in $s_{0}$ and $tn$ decomposes into a primitive task network $(T^{*},{\prec^{+}}^{*},\alpha^{*})$ with $T^{\prime}\subseteq T^{*}$, for each $(t,t^{\prime})\in{\prec^{+}}^{*}$ with $t^{\prime}\in T^{\prime}$ we have that $t\in T^{\prime}$ and $(t,t^{\prime})\in{\prec^{+}}^{\prime}$, and $\alpha^{\prime}(t)=\alpha^{*}(t)$ for all $t\in T^{\prime}$. The literature usually considers the case where $tn$ can be decomposed into $tn^{\prime}$, that is, $T^{*}=T^{\prime}$, in which we call $tn^{\prime}$ a full solution. Intuitively, non-full solutions respect $\delta$ and $\prec^{+}$, but do not need to execute all tasks. We now define our problems of interest.
@@
Plan Verification Input: A task network $tn$ in a domain $D$, an initial state $s_{0}\subseteq F$, and a plan $p$ over $A$. Question: Is there a full solution to $(D,tn,s_{0})$ with a linearization $\bar{t}$ such that $\alpha(\bar{t})=p$?